OpenStudy (anonymous):

how do i factor 4k^3 +14k^2+ 14k + 4

OpenStudy (anonymous):

any ideas? I didn't know if i should separate them into two problems

OpenStudy (anonymous):

Hi :-) make 14k common between middle terms and 4 between 4k^3 and 4

OpenStudy (anonymous):

so would it become 4(k^3) + 14k(k+1)

OpenStudy (anonymous):

you will come up with: \[4(k^3+1)+14k(k+1)\] then make use of this identity:\[a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

I don't know how to use that identity

OpenStudy (anonymous):

ok, I'll tell you :-) let \(a=k\) and \(b=1\) so you have:\[k^3+1=(k+1)(k^2-k+1)\]am i clear?

OpenStudy (anonymous):

ok I am with you so far, what happens to the 4?

OpenStudy (anonymous):

well, let's see what we have till here\[4(k+1)(k^2-k+1)+14k(k+1)\]now make 2(k+1) common, what will you get?

OpenStudy (anonymous):

2(k+1)(k^2-k+1)(14k+4)

OpenStudy (anonymous):

there is a mistake, give it a try again

OpenStudy (anonymous):

(k+1)^2 (K^2-k+1) (14k+4)

OpenStudy (anonymous):

Is it right this time?

OpenStudy (anonymous):

sry, give me a minute, I'm coming

OpenStudy (anonymous):

oh my bad

OpenStudy (anonymous):

@hartnn plz help middi :-) i have to go

hartnn (hartnn):

sure

hartnn (hartnn):

\(4(k+1)(k^2-k+1)+14k(k+1)\) factoring 2 and k+1 from the 1st term, what remains is 2 and k^2 -k+1 right ? from the 2nd term, what remains is 7 and k

hartnn (hartnn):

got this ? \( 4(k+1)(k^2-k+1)+14k(k+1) \\ 2(k+1) [2 (k^2 -k+1) +7k] \)

hartnn (hartnn):

you can simplify the terms inside [...] bracket

OpenStudy (anonymous):

so would i start by applying 2 to what is inside the parenthesis

hartnn (hartnn):

thats correct, do that

OpenStudy (anonymous):

2k^2 +5k +1

hartnn (hartnn):

just one error in that : 2k^2 -2k +2 +7k = ...

OpenStudy (anonymous):

whoops 2k^2 +5k + 2

hartnn (hartnn):

yes now thats easily factor-able can you tell me 2 numbers with product = 2*2 and sum = 5 ?

OpenStudy (anonymous):

(2k +1)(k+2)

hartnn (hartnn):

correct! now put everything together :)

OpenStudy (anonymous):

2(k+1)(2k +1)(k+2)

hartnn (hartnn):

2 (k+1) which we got previously so, \(\large 2 (2k+1)(k+1)(k+2)\)

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

thank you :)

hartnn (hartnn):

welcome ^_^